OpenStudy (jennychan12):

What is this saying? Find the area between the curves y = x/2 -1 and y = cos x., bounded on the left by the y axis.

OpenStudy (jennychan12):

\[y = \frac{ 1 }{ 2 }x-1\] \[y = \cos x\] looks something like this: